OpenStudy (anonymous):

Express the complex number in trigonometric form. -2

jimthompson5910 (jim_thompson5910):

hint: -2 = -2 + 0i

OpenStudy (anonymous):

2(cos 0° + i sin 0°)?

jimthompson5910 (jim_thompson5910):

any number in a+bi form can be converted to trigonometric form z = r*[ cos(theta) + i*sin(theta) ] where r = sqrt(a^2 + b^2) theta = arctan(b/a)

jimthompson5910 (jim_thompson5910):

since -2+0i has a negative component, this means that you'll have to add on 180 degrees to theta (to make sure you're in the right region)

jimthompson5910 (jim_thompson5910):

2(cos 0° + i sin 0°) is incorrect

OpenStudy (anonymous):

so then it would be 2(cos 180° + i sin 180°)?

jimthompson5910 (jim_thompson5910):

to find r, use this formula r = sqrt(a^2 + b^2)

jimthompson5910 (jim_thompson5910):

in this case, a = -2, b = 0

OpenStudy (anonymous):

I'm lost now.

jimthompson5910 (jim_thompson5910):

-2 = -2+0i if you were to plot this point on the xy axis (x is the real number line, y is the imaginary number line), then you'd be plotting the point (-2,0)

jimthompson5910 (jim_thompson5910):

oh wait, I misread and didn't see you posted 2(cos 180° + i sin 180°) 2(cos 180° + i sin 180°) is correct

OpenStudy (anonymous):

thanks